# Let’s Measure Information Step-by-Step: LLM-Based Evaluation Beyond Vibes

### v1

@@ A.3.1 Stage 1: Reading and Note-Taking
A.3.2 Stage 2: Review Generation
@@ A.4.1 TVD-MI Critic
A.4.2 LLM Judge (Without Context)
@@
Appendix B Summarization Experiment Details
@@ B.2 Experimental Configuration
B.3 Agent Summarization Strategies
@@ 29 Minimal Detail Avoid specific details from the article Low Effort
B.4 Example Strategy Implementations
B.4.1 Faithful Summarization
@@
B.4.2 Fact Manipulation (Strategic)
@@
B.4.3 Conspiracy Theory (Strategic)
@@
B.4.4 Ultra Concise (Low Effort)
@@ B.5 Evaluation Protocol
B.6 Statistical Analysis
Despite explicit instructions for 3-4 bullet points, generated summaries averaged 105.7 words for the Faithful condition which is 2.4× longer than reference summaries. This systematic verbosity across conditions (except Ultra Concise at 15.2 words) reveals an inherent bias in the model’s summarization behavior, creating additional challenges for evaluation mechanisms to distinguish quality from length effects. Importantly, our mechanisms show weak correlations with length (|r| < 0.4), confirming they capture information-theoretic properties beyond simple verbosity.
Appendix C Machine Translation Experiment Details
@@ C.2 Experimental Setup
C.3 Agent Translation Strategies
@@ C.4 Categorization Rationale
C.5 Example Strategy Implementations
C.5.1 Faithful Translation (Original)
@@
C.5.2 Strategic Manipulation (Misleading)
@@
C.5.3 Style Variation (Poetic)
@@ C.6 Evaluation Protocol
Appendix D Proofs
D.1 Binary Case: $\kappa$ as Normalized TVD-MI and General Lower Bound
@@
Writing the 2×2 contingency table with cells $P_{00},P_{01},P_{10},P_{11}$, one shows:
$\mathrm{TVD}\bigl{(}P_{X,Y},P_{X}P_{Y}\bigr{)}=\tfrac{1}{2}\sum_{i,j\in\{0,1\}}\bigl{|}P_{ij}-P_{X}(i)P_{Y}(j)\bigr{|}=p_{o}-p_{e}.$
@@
$\boxed{\kappa=\frac{\mathrm{TVD}}{1-p_{e}}\quad\Longleftrightarrow\quad\mathrm{TVD}=(1-p_{e})\,\kappa.}$
@@
$\mathrm{TVD}\bigl{(}P,P_{X}P_{Y}\bigr{)}=\tfrac{1}{2}\sum_{i,j}\bigl{|}p_{ij}-p_{i\cdot}p_{\cdot j}\bigr{|}\geq\tfrac{1}{2}\sum_{i}\bigl{(}p_{ii}-p_{i\cdot}p_{\cdot i}\bigr{)}=\tfrac{1}{2}(p_{o}-p_{e})=\tfrac{1}{2}\,\kappa\,(1-p_{e}).$
@@
This shows that in the general (multi-category) case, Cohen’s $\kappa$ provides a lower bound (up to normalization) on the total variation distance between the joint and the product of marginals, justifying TVD-MI as a natural extension of inter-rater reliability measures.
D.2 Proof of Theorem 3.3 and 3.5
Before we prove our result we first prove the following lemma which establishes when we can maximize $f$-mutual information.
Lemma D.1.
@@
Consider a distribution $p_{X,Y}$ and $N\geq 2$. We denote by $I_{f}^{*}(N)$ the maximum attainable mutual information with $N$ elements in the support. If the support of $p_{X,Y}$ has fewer than $2kN^{2}$ elements then $I_{f}(X;Y)<I_{f}^{*}(2kN^{2})$ and by the premise of the theorem we have that, with probability at least $1-\delta$ over the draw of $S$, $B(\mathcal{T}(S),\delta)\leq I_{f}(X;Y)$ so the theorem follows.
@@
We will let $\mathrm{Small}(S)$ denote the event that $B(\mathcal{T}(S),\delta)\leq I_{f}^{*}(2kN^{2})$ and let $\mathrm{Pure}(S)$ abbreviate the event that no element $z_{i}$ for $i>kN^{2}$ occurs twice in the sample. Since $\tilde{p}_{X,Y}$ has a support of size $2kN^{2}$ we have
$I_{f}(X;Y)\leq I_{f}^{*}(2kN^{2})=\frac{1}{2kN^{2}}f(2kN^{2})+\left(1-\frac{1}{2kN^{2}}\right)f(0).$
which follows from Lemma D.1. Applying our hypothesis to $\tilde{p}_{X,Y}$ gives
$\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Small}(S))\geq 1-\delta.$
For a type $\mathcal{T}(S)$ let $\Pr{S\sim P^{N}}(T)$ denote the probability over drawing $S\sim P^{N}$ that $\mathcal{T}(S)=T$. We now have
$\Pr_{S\sim p_{X,Y}^{N}}(T|\mathrm{Pure}(S))=\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(T|\mathrm{Pure}(S)).$
This gives the following
$\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Small}(S))$ $\geq\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Pure}(S)\wedge\mathrm{Small}(S))$ (8)
$=\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Pure}(S))\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Small}(S)|\mathrm{Pure}(S))$ (9)
$=\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Pure}(S))\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Small}(S)|\mathrm{Pure}(S))$ (10)
$\geq\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Pure}(S))\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Pure}(S)\wedge\mathrm{Small}(S)).$ (11)
For $i>kN^{2}$ we have $\tilde{p}_{X,Y}(z_{i})\leq 1/(kN^{2})$ and then we consider the complement event $\lnot\text{Pure}(S)$ which is the event that some tail element appears at least twice. There are $\binom{N}{2}$ different valid index combinations so we get the following from a union bound
$1-\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\text{Pure}(S))$ $=\binom{N}{2}\cfrac{1}{kN^{2}}=\cfrac{N(N-1)}{2kN^{2}}\leq\frac{1}{2k}$ (12)
$\Rightarrow\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\text{Pure}(S))$ $\geq 1-\frac{1}{2k}.$ (13)
Applying the union bound to the previous inequality and $\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Small}(S))$ gives
$\Pr_{S\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Pure}(S)\wedge\mathrm{Small}(S))\geq 1-\delta-\frac{1}{2k}.$
We also know $p_{X,Y}(z_{kN^{2}+i})\leq\frac{1}{kN^{2}}$ or else $\sum_{i\leq kN^{2}}p_{X,Y}(z_{i})\geq 1$. So by a derivation similar to that above we get
$\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Pure}(S))\geq 1-\frac{1}{2k}.$
Combining this with another union bound gives
$\Pr_{S\sim p_{X,Y}^{N}}(\mathrm{Small}(S))\geq 1-\delta-\frac{1}{k}$

### v2

@@ C.3.1 Stage 1: Reading and Note-Taking
C.3.2 Stage 2: Review Generation
@@ C.4.1 TVD-MI Critic
C.4.2 LLM Judge (Without Context)
@@
Appendix D Summarization Experiment Details
@@ D.2 Experimental Configuration
D.3 Agent Summarization Strategies
@@ 29 Minimal Detail Avoid specific details from the article Low Effort
D.4 Example Strategy Implementations
D.4.1 Faithful Summarization
@@
D.4.2 Fact Manipulation (Strategic)
@@
D.4.3 Conspiracy Theory (Strategic)
@@
D.4.4 Ultra Concise (Low Effort)
@@ D.5 Evaluation Protocol
D.6 Statistical Analysis
Despite explicit instructions for 3-4 bullet points, generated summaries averaged 105.7 words for the Faithful condition which is 2.4× longer than reference summaries. This systematic verbosity across conditions (except Ultra Concise at 15.2 words) reveals an inherent bias in the model’s summarization behavior, creating additional challenges for evaluation mechanisms to distinguish quality from length effects. Additionally, our mechanisms show weak correlations with length (|r| < 0.4), confirming they capture information-theoretic properties beyond simple verbosity.
Appendix E Machine Translation Experiment Details
@@ E.2 Experimental Setup
E.3 Agent Translation Strategies
@@ E.4 Categorization Rationale
E.5 Example Strategy Implementations
E.5.1 Faithful Translation (Original)
@@
E.5.2 Strategic Manipulation (Misleading)
@@
E.5.3 Style Variation (Poetic)
@@ E.6 Evaluation Protocol
Appendix F Proofs
F.1 $\text{Cohen's }\kappa$ as Normalized TVD-MI and General Relationships
@@
Writing the $2\times 2$ contingency table with cells $P_{00},P_{01},P_{10},P_{11}$, one has:
$\mathrm{TVD}\bigl{(}P_{X,Y},P_{X}P_{Y}\bigr{)}=\tfrac{1}{2}\sum_{i,j\in\{0,1\}}\bigl{|}P_{ij}-P_{X}(i)P_{Y}(j)\bigr{|}\;\;\geq\;\;\tfrac{1}{2}\,(p_{o}-p_{e}).$
@@
$\boxed{\;|\kappa|\;\leq\;\frac{2\,\mathrm{TVD}}{1-p_{e}}\quad\Longleftrightarrow\quad\mathrm{TVD}\;\geq\;\tfrac{1}{2}\,(1-p_{e})\,|\kappa|\;.}$
@@
$\mathrm{TVD}\bigl{(}P,P_{X}P_{Y}\bigr{)}=\tfrac{1}{2}\sum_{i,j}\bigl{|}p_{ij}-p_{i}\cdot p_{j}\bigr{|}\geq\tfrac{1}{2}\sum_{i}\bigl{|}p_{ii}-p_{i}\cdot p_{i}\bigr{|}\geq\tfrac{1}{2}\,\bigl{(}p_{o}-p_{e}\bigr{)}=\tfrac{1}{2}\,\kappa\,(1-p_{e}).$
@@
This shows that in the general (multi-category) case, Cohen’s $\kappa$ provides a lower bound (up to normalization) on the total variation distance between the joint and the product of marginals, justifying TVD-MI as a natural extension of inter-rater reliability measures. In high-dimensional settings, such as text, we expect $p_{e}\sim 0$, allowing $\kappa\lessapprox 2\,\mathrm{TVD}$.
Unification with AUC and Informativeness
Building on Powers [Powers, 2012], we can show that for binary decisions:
  1. 1.
    TVD-MI and Informativeness: For balanced prevalence, TVD-MI = (TPR + TNR - 1)/2 = Youden’s J/2
  2. 2.
    Informativeness and AUC: Youden’s J = 2(AUC - 0.5) when the ROC curve is symmetric
  3. 3.
    $\kappa$ and Informativeness: $\kappa\approx$ Informativeness when chance agreement is low
This trinity of relationships explains our empirical findings:
  1. 1.
    Why TVD-MI successfully produces item-level AUC scores (Table 3)
  2. 2.
    Why our mechanisms correlate with quality metrics where ground truth exists
  3. 3.
    Why optimizing for gaming-resistance (via TVD-MI) simultaneously optimizes for discrimination (AUC)
A key insight from Powers [Powers, 2012] is that these measures all capture the same underlying concept. This is the degree to which classifications contain information beyond chance, but with different normalizations suited to different contexts.
F.2 Proof of Theorem 3.3
Before we present our result we first show the following lemma which establishes when we can maximize $f$-mutual information.
Lemma F.1.
@@
Consider a distribution $p_{X,Y}$ and $N\geq 2$. We denote by $I_{f}^{*}(N)$ the maximum attainable mutual information with $N$ elements in the support. If the support of $p_{X,Y}$ has fewer than $2kN^{2}$ elements then $I_{f}(X;Y)<I_{f}^{*}(2kN^{2})$ and by the premise of the theorem we have that, with probability at least $1-\delta$ over the draw of $S^{(N)}$, $B(\mathcal{T}(S^{(N)}),\delta)\leq I_{f}(X;Y)$ so the theorem follows.
@@
We will let $\mathrm{Small}(S^{(N)})$ denote the event that $B(\mathcal{T}(S^{(N)}),\delta)\leq I_{f}^{*}(2kN^{2})$ and let $\mathrm{Pure}(S^{(N)})$ abbreviate the event that no element $z_{i}$ for $i>kN^{2}$ occurs twice in the sample. Since $\tilde{p}_{X,Y}$ has a support of size $2kN^{2}$ we have
$I_{f}(X;Y)\leq I_{f}^{*}(2kN^{2})=\frac{1}{2kN^{2}}f(2kN^{2})+\left(1-\frac{1}{2kN^{2}}\right)f(0),$
which follows from Lemma F.1. Applying our hypothesis to $\tilde{p}_{X,Y}$ gives
$\Pr_{S^{(N)}\sim\tilde{p}_{X,Y}^{N}}(\mathrm{Small}(S^{(N)}))\geq 1-\delta.$
Couple $S^{(N)}\sim p_{X,Y}^{N}$ and $\tilde{S}^{(N)}\sim\tilde{p}_{X,Y}^{N}$ by using the same draws on the head $\{z_{1},\dots,z_{kN^{2}}\}$ and drawing tail samples independently according to their respective tail distributions. On the event $\mathrm{Pure}(S^{(N)})\wedge\mathrm{Pure}(\tilde{S}^{(N)})$ we have $\mathcal{T}(S^{(N)})=\mathcal{T}(\tilde{S}^{(N)})$, hence
$\Pr_{p_{X,Y}^{N}}(\neg\mathrm{Small})\leq\Pr\bigl{(}\mathcal{T}(S^{(N)})\neq\mathcal{T}(\tilde{S}^{(N)})\bigr{)}+\Pr_{\tilde{p}_{X,Y}^{N}}(\neg\mathrm{Small})\leq\Pr_{p_{X,Y}^{N}}(\neg\mathrm{Pure})+\Pr_{\tilde{p}_{X,Y}^{N}}(\neg\mathrm{Pure})+\delta.$
For $i>kN^{2}$ we have $\tilde{p}_{X,Y}(z_{i})\leq 1/(kN^{2})$. Consider the complement event $\lnot\mathrm{Pure}(S^{(N)})$ that some tail element appears at least twice. By a union bound over the $\binom{N}{2}$ index pairs and using $\sum_{i>kN^{2}}q_{i}^{2}\leq\max_{i}q_{i}\cdot\sum_{i>kN^{2}}q_{i}\leq 1/(kN^{2})$ for the tail distribution $(q_{i})$, we obtain
$\Pr_{S^{(N)}\sim\tilde{p}_{X,Y}^{N}}(\lnot\mathrm{Pure}(S^{(N)}))$ $\leq\binom{N}{2}\cdot\frac{1}{kN^{2}}=\frac{N(N-1)}{2kN^{2}}\leq\frac{1}{2k},$ (6)
$\Pr_{S^{(N)}\sim p_{X,Y}^{N}}(\lnot\mathrm{Pure}(S^{(N)}))$ $\leq\binom{N}{2}\cdot\frac{1}{kN^{2}}\leq\frac{1}{2k},$ (7)
where for $p_{X,Y}$ we also used $p_{X,Y}(z_{kN^{2}+i})\leq 1/(kN^{2})$ (else $\sum_{i\leq kN^{2}}p_{X,Y}(z_{i})\geq 1$).
Plugging these bounds into $(\star)$ yields
$\Pr_{p_{X,Y}^{N}}(\neg\mathrm{Small})\leq\delta+\frac{1}{2k}+\frac{1}{2k}=\delta+\frac{1}{k},$
i.e.
$\Pr_{S^{(N)}\sim p_{X,Y}^{N}}(\mathrm{Small}(S^{(N)}))\geq 1-\delta-\frac{1}{k},$